OpenStudy (anonymous):

Solve for y, then find the value of y when given x = –2.

OpenStudy (anonymous):

what is the equation?

OpenStudy (anonymous):

Where is the rest of the problem?

OpenStudy (anonymous):

6x = 7 – 4y (1 point)

OpenStudy (anonymous):

-12=7-4y -19=4y y=4 3/4

OpenStudy (anonymous):

wait y =-4 3/4

OpenStudy (anonymous):

Plug in x, to the second equation. 6(-2) = 7-4y Simplify -8=7-4y Simplify -15=-4y Divide by -4 to both sides y is 3 3/4.

OpenStudy (anonymous):

Okay, Thanks A Bunch! 

OpenStudy (anonymous):

Ahhh I messed up. Ignore me.

OpenStudy (anonymous):

grammar you messed up

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

you forgot to multiply XD

OpenStudy (anonymous):

Shhh you didn't see anything

OpenStudy (anonymous):

haha, its okay! :3